OpenStudy (anonymous):

A hardware supplier manufactures three kinds of clamps, types A, B, and C. Production restrictions force it to make ten more C clamps than the total of the other types and three times as many B clamps as type A clamps. The shop must produce 490 clamps per day. How many of each type are made per day??? Help me please???a minute ago

OpenStudy (dumbcow):

(i) a+b+c = 490 (ii) c = a+b+10 (iii) b = 3a substitute c from (ii) into (i) and b from (iii) into (i) a + 3a + (a+3a+10) =490 8a = 490-10 a =60 b=3*60 = 180 c = 60+180+10 = 250

OpenStudy (anonymous):

thx